# Trilinos: Enabling Scientific Computing Across Diverse Hardware Architectures at Scale

### v1

@@ Abstract.
copyright: acmcopyrightjournalyear: 2025doi: XXXXXXX.XXXXXXXccs: Mathematics of computing Solversccs: Mathematics of computing Mathematical software performanceccs: Computing methodologies Massively parallel algorithmsccs: Software and its engineering Software libraries and repositoriesccs: Software and its engineering Software performance
@@ 1. Introduction
Trilinos was originally conceived as a framework of three packages for distributed memory systems. The original Trilinos publication (Heroux et al., 2005) described the motivation, philosophy, and capabilities of Trilinos at that time. Seven years later, a two-part collection of 16 papers provided a new overview of Trilinos and many of its packages (Heroux and Willenbring, 2012; Heroux and Rouson, 2012a, b; Pawlowski et al., 2012a; Oldfield et al., 2012; Bochev et al., 2012; Boman et al., 2012; Baker and Heroux, 2012; Edwards et al., 2012; Pawlowski et al., 2012b; Spotz, 2012; Long et al., 2012; Morris et al., 2012; Howle et al., 2012; Bavier et al., 2012; Gaidamour et al., 2012) and described the intervening expansion of Trilinos’ capabilities and strategic goals for Trilinos. Trilinos today hews to the overall vision from two decades ago, but has evolved alongside changes in programming models, application needs, hardware architectures, and algorithms. Trilinos has grown from a library of three packages to more than fifty packages with functionality and features supporting a wide range of applications.
For much of its history, packages have relied heavily on the linear algebra classes provided by Epetra. Epetra does not support larger problems (2B+ unknowns), nor does it allow for accelerator offload. Since the packages of the old Epetra stack are scheduled to be archived and removed from the main Trilinos repository on GitHub by the end of 2025, this paper describes only the modern Trilinos software stack, which builds upon Tpetra and the Kokkos ecosystem to achieve performance portability across hardware architectures.
This article attempts to capture a snapshot the current state of Trilinos, as an update to the articles published thirteen and twenty years ago. Therefore it focuses on the major developments within Trilinos in the last decade as well as new features and functionality that have been added to advance scientific and engineering applications. It will give only an overview of the features, and we refer to the extensive list of references for the details of these features. We are also cognizant of the fact that this article describes software that is being actively developed and constantly evolving. Hence, we will focus on the high-level features and concepts that we expect to remain stable for several years. We also briefly touch upon the Trilinos community and software engineering issues with respect to Trilinos.
@@ 1.1. Modern Trilinos: Performance Portability through Kokkos
[…]
@@
Kokkos enables Trilinos developers to write single source implementations of their packages that perform well on all major HPC hardware architectures. Some of its design principles are also reflected in Trilinos API designs. In particular, the use of execution and memory spaces as parameters of data structures and algorithms is a common thread throughout Trilinos. This design characteristic enables Trilinos users to manage the heterogeneity of modern architectures and its impact on Trilinos data structures.
@@ 1.2. Trilinos Functionality
While most packages are developed natively in the Trilinos repository, some packages are developed externally and snapshotted into Trilinos. These are (in alphabetical order): Compadre [^4] https://github.com/sandialabs/compadre, Kokkos and Kokkos Kernels, Krino, SEACAS [^5] https://github.com/sandialabs/seacas, Sierra ToolKit (STK), and TriBITS [^6] https://github.com/TriBITSPub/TriBITS. Some reasons for snapshotting external packages into Trilinos include: external packages and repositories facilitate easy and independent usage by other software projects without having to pull in Trilinos; snapshotting provides an integration process for external packages into Trilinos, ensuring continual compatibility between all packages; and snapshotting allows easy integration into Trilinos-based application codes, alleviating users from managing these dependencies themselves. We refer to the Trilinos website (https://trilinos.github.io) for a brief description of these packages.
In the following sections of this paper, we group the Trilinos packages that share common objectives (e.g., solving linear systems) together into Trilinos’ five product areas: Core, Linear Solvers and Preconditioners, Nonlinear Solvers and Analysis Tools, Discretization Tools, and Framework Infrastructure. Trilinos’ packages and their assignment to a product area are illustrated in Figure 1.
@@
These packages provide high-level algorithms for computational simulations and design. Capabilities include solvers for nonlinear equations, parameter continuation, bifurcation tracking, optimization, and uncertainty quantification. Trilinos also provides lower-level utility packages to evaluate quantities of interest required by the analysis algorithms. Capabilities include automatic differentiation technology to evaluate derivatives and embedded ensemble propagation for uncertainty quantification. These packages and their capabilities will be discussed further in Section 4.
@@
This collection of packages provides functionality for the discretization of differential equations. In particular, it supports mesh-free and mesh-based spatial discretizations, with a focus on high-order finite elements, and time integration. Discretization tools also include cross-cutting utilities for algorithmic differentiation and for managing directed acyclic graphs of evaluation kernels. These capabilities are described further in Section 5 in detail.
@@ 2.1. Performance Portability: Kokkos Core
Kokkos Core [^7] For historical reasons the package name is “Kokkos” since it precedes the creation of other Kokkos subprojects such as Kokkos Kernels. is a programming model designed for performance portability across hardware architectures such as CPUs and GPUs. Two primary insights into achieving performance portability shaped Kokkos’ development: i) abstractions for both parallel execution and data structures are required, and ii) the fundamental programming model must be descriptive rather than prescriptive.
The Kokkos programming model includes execution spaces to specify where an algorithm should run, execution policies to indicate how the algorithm should be run in terms of hardware resources, and execution patterns to designate whether the for, reduce, or scan pattern will be used by the algorithm. The data structure abstractions are built around the concepts of memory spaces to indicate where memory is allocated, memory layouts to specify how data is laid out in memory, and memory traits that enable customizations of how data is accessed. The primary data structure incorporating these concepts is Kokkos::View, a multidimensional array abstraction that inspired the ISO C++23 mdspan class [^8] https://www.open-std.org/jtc1/sc22/wg21/docs/papers/2022/p0009r18.html. Detailed descriptions of Kokkos Core can be found in (Edwards et al., 2014) and (Trott et al., 2022).
@@ 2.2. Performance Portable Kernels: Kokkos Kernels
[…]
@@
The capabilities provided by Kokkos Kernels can be divided into four major categories: 1. BLAS algorithms, 2. sparse linear algebra and preconditioners, 3. graph algorithms, and 4. batched dense and sparse linear algebra (Liegeois et al., 2023). The main points of integration of Kokkos Kernels in Trilinos are Tpetra for the dense and sparse linear algebra capabilities, Ifpack2 for the preconditioners and batched algorithms, and the multigrid package MueLu which relies directly and indirectly on these general capabilities as well as more specialized algorithms such as graph coloring/coarsening (Kelley and Rajamanickam, 2022) and fused Jacobi-SpGEMM (generalized sparse matrix-matrix multiply) kernels.
@@ 2.3. Tools: Teuchos
Teuchos provides a suite of common tools for many Trilinos packages. These tools include memory management classes (Bartlett, 2010) such as smart pointers and arrays, parameter lists for communicating hierarchical lists of parameters between library or application layers, templated wrappers for BLAS and LAPACK, XML parsers, MPI, and other utilities. They provide a unified “look and feel” across Trilinos packages and help avoid common programming mistakes.
@@ 2.4. Distributed-Memory Linear Algebra: Tpetra
Tpetra (Baker and Heroux, 2012; Hoemmen, 2015) provides the distributed-memory infrastructure for linear algebra objects such as sparse graphs, sparse matrices, and dense vectors. These linear algebra objects are implemented using MPI-based abstractions for the inter-process communication, and Kokkos-based data structures for the portions of the object owned by each process (MPI rank). Kokkos Kernels is used for performing local, parallel computations within each MPI rank.
The objects provided are templated on scalar type (e.g., double, float, or a Sacado or Stokhos type), local index type, global index type, and Kokkos device (pair of execution and memory spaces). The distribution of the linear algebra objects is described by a Map data structure, which defines which elements are present on which ranks. Here “element” could be a row or column of a sparse matrix or an entry of a vector. Global indices number elements across an entire distributed object, while local indices number elements only within a rank. Since memory capacity limits the sizes of local objects, local indices can be safely represented with a more compact type than global indices (e.g., 32-bit int instead of 64-bit long).
@@
  • (Multi)Vectors: storage of dense vectors or collections of vectors (multivectors) and associated BLAS-1 like kernels (e.g., dot products, norms, scaling, vector addition, pointwise vector multiplication) as well as tall skinny QR (TSQR) factorization for multivectors.
  • CrsGraph a sparse graph in compressed row storage (CRS) format. Graphs also include import/export objects for use in halo/boundary exchanges associated with the graph.
  • CrsMatrix and BlockCrsMatrix are sparse matrices in CRS and block compressed sparse row (BSR) formats. Associated kernels include sparse matrix-vector product (SPMV), sparse matrix-matrix multiplication (SPGEMM) and triple-product, sparse matrix-matrix addition, sparse matrix transpose, diagonal extraction, Frobenius norm calculation, and row/column scaling.
@@ 2.5. Abstract Numerical Algorithms: Thyra
The Thyra package provides abstractions for implementing high-level abstract numerical algorithms (ANAs). Thyra is used by other Trilinos packages for implementing ANAs such as solvers for linear and nonlinear equations, bifurcation and stability analysis, optimization, and uncertainty quantification. The abstractions hide concrete linear algebra and solver implementations so that the algorithms can be written once and reused across many application codes. For example, one could write a Krylov solver and use it with multiple concrete linear algebra implementations such as Tpetra in Trilinos or the vector/matrix objects in PETSc (Balay et al., 2021).
@@
  • Operator Solve Interfaces: The second set of abstractions supplies abstract linear solvers that can be attached to operators. The interfaces include base classes for preconditioners, preconditioner factories, linear operators with a solver, and factories for linear operators with a solver. The factories build and update the preconditioners and linear solvers as needed. The solver interfaces allow one to wrap any concrete linear solver or preconditioner for use with the ANAs. Trilinos provides a separate library named Stratimikos that wraps all of the Trilinos preconditioners and linear solvers in Thyra abstractions. Run-time configuration of Stratimikos is driven by Teuchos::ParameterList options. Details on Stratimikos are given in Section 3.8.
  • Nonlinear Interfaces: The final set of abstractions supports nonlinear operations. Thyra provides an interface that wraps nonlinear solvers. Implicit time integration and optimization solvers typically require nonlinear solves and can abstract away the concrete implementation via Thyra. The NOX package in Trilinos (described in Section 4.1) provides a concrete implementation of this interface. The final interface in Thyra is for querying a physics application for common objects required by high-level Newton-based solvers. The Thyra::ModelEvaluator interface allows codes to ask the application to evaluate residuals, Jacobians, parametric sensitivities, Hessians, and post-processed quantities/derivatives of interest. The interface is designed to be stateless (i.e., consecutive calls with the same inputs should produce the same outputs). This design was chosen to avoid pitfalls in complex application codes interacting with general nonlinear algorithms. The design does not preclude its use in stateful applications that have a “memory” between timesteps, but rather forces the explicit use of the stateful information in the interface.

### v2

@@ Abstract.
copyright: acmcopyrightjournalyear: 2025doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8
@@ 1. Introduction
Trilinos was originally conceived as a framework of three packages for distributed memory systems. The original Trilinos publication (Heroux et al., 2005) described the motivation, philosophy, and capabilities of Trilinos at that time. Seven years later, a two-part collection of 16 papers provided a new overview of Trilinos and many of its packages (Heroux and Willenbring, 2012; Heroux and Rouson, 2012a, b; Pawlowski et al., 2012b; Oldfield et al., 2012; Bochev et al., 2012; Boman et al., 2012; Baker and Heroux, 2012; Edwards et al., 2012; Pawlowski et al., 2012a; Spotz, 2012; Long et al., 2012; Morris et al., 2012; Howle et al., 2012; Bavier et al., 2012; Gaidamour et al., 2012) and described the intervening expansion of Trilinos’ capabilities and strategic goals for Trilinos. Trilinos today hews to the overall vision from two decades ago, but has evolved alongside changes in programming models, application needs, hardware architectures, and algorithms. Trilinos has grown from a library of three packages to more than 35 packages with functionality and features supporting a wide range of applications.
For much of its history, packages have relied heavily on the linear algebra classes provided by Epetra. Epetra did not support larger problems (2B+ unknowns), nor did it allow for accelerator offload. Since the packages of the old Epetra stack were archived and removed from the main Trilinos repository on GitHub by the end of 2025, this paper describes only the modern Trilinos software stack, which builds upon Tpetra and the Kokkos ecosystem to achieve performance portability across hardware architectures.
This article attempts to capture a snapshot the current state of Trilinos, as an update to the articles published 14 and 21 years ago. Therefore it focuses on the major developments within Trilinos in the last decade as well as new features and functionality that have been added to advance scientific and engineering applications. It will give only an overview of the features, and we refer to the extensive list of references for the details of these features. We are also cognizant of the fact that this article describes software that is being actively developed and constantly evolving. Hence, we will focus on the high-level features and concepts that we expect to remain stable for several years. We also briefly touch upon the Trilinos community and software engineering issues with respect to Trilinos.
@@ 1.1. Modern Trilinos: Performance Portability through Kokkos
A key goal of modern Trilinos is to offer a performance-portable collection of reusable scientific libraries, allowing users to develop applications that achieve high efficiency across all modern high performance computing (HPC) hardware architectures. This goal emerged when the HPC architecture landscape started diversifying with the introduction of GPU acceleration for scientific software. Today, nine of the top ten systems in the Top500111https://top500.org, 66th edition from Nov. 17, 2025 use GPUs, with only a single system using CPUs only. A challenge in this shift is the diversification of vendor-provided programming models: the aforementioned systems have four different preferred programming models: CUDA, OpenACC, HIP and SYCL for the GPU-based systems and OpenMP for the CPU system.
@@
Kokkos enables Trilinos developers to write single source implementations of their packages that perform well on all major HPC hardware architectures. Some of its design principles are also reflected in Trilinos API designs. In particular, the use of execution spaces (where does code run) and memory spaces (where does data reside) as parameters of data structures and algorithms is a common thread throughout Trilinos. This design characteristic enables Trilinos users to manage the heterogeneity of modern architectures and its impact on Trilinos data structures.
@@ 1.2. Trilinos Functionality
While most packages are developed natively in the Trilinos repository, some packages are developed externally and “snapshotted” into Trilinos, i.e., incorporated as fixed copies of their source code at a specific point in time. These are (in alphabetical order): Compadre [^4] https://github.com/sandialabs/compadre, Kokkos and Kokkos Kernels, Krino, ROL [^5] https://github.com/sandialabs/rol, SEACAS [^6] https://github.com/sandialabs/seacas, Sierra ToolKit (STK), and TriBITS [^7] https://github.com/TriBITSPub/TriBITS. Some reasons for snapshotting external packages into Trilinos include: external packages and repositories facilitate easy and independent usage by other software projects without having to pull in Trilinos; snapshotting provides an integration process for external packages into Trilinos, ensuring continual compatibility between all packages; and snapshotting allows easy integration into Trilinos-based application codes, alleviating users from managing these dependencies themselves. We refer to the Trilinos website (https://trilinos.github.io) for a brief description of these packages.
In the following sections of this paper, we group the Trilinos packages that share common objectives (e.g., solving linear systems) together into Trilinos’ five product areas: Core, Linear Solvers and Preconditioners, Nonlinear Solvers and Analysis Tools, Discretization Tools, and Framework. Trilinos’ packages and their assignment to a product area are illustrated in Figure 1.
@@
These packages provide algorithms for computational simulations and design. Capabilities include solvers for nonlinear equations, parameter continuation, bifurcation tracking, optimization, and uncertainty quantification. Trilinos also provides utility packages to evaluate quantities of interest required by the analysis algorithms. Additional capabilities include automatic differentiation technology to evaluate derivatives and embedded ensemble propagation for uncertainty quantification. These packages and their capabilities will be discussed further in Section 4.
@@
This collection of packages provides functionality for the discretization of differential equations. In particular, it supports mesh-free and mesh-based spatial discretizations, with a focus on high-order finite elements, and time integration. Discretization tools also include support for algorithmic differentiation and for managing directed acyclic graphs of evaluation kernels. These capabilities are described further in Section 5 in detail.
@@ 1.3. Outline of this manuscript
In comparison to the first Trilinos publication from 2005 (Heroux et al., 2005) and its updates from 2012 (Heroux and Willenbring, 2012; Heroux and Rouson, 2012a, b; Pawlowski et al., 2012b; Oldfield et al., 2012; Bochev et al., 2012; Boman et al., 2012; Baker and Heroux, 2012; Edwards et al., 2012; Pawlowski et al., 2012a; Spotz, 2012; Long et al., 2012; Morris et al., 2012; Howle et al., 2012; Bavier et al., 2012; Gaidamour et al., 2012), the present manuscript not only provides updates on newly added capabilities, but more importantly reflects the overhaul of Trilinos towards performance portability through the full integration of the Kokkos ecosystem (Edwards et al., 2014; Trott et al., 2021, 2022). Many packages described in the original Trilinos publication (Heroux et al., 2005) have been replaced by Kokkos-capable successor implementations, among them (in alphabetical order) Amesos2 replacing Amesos, Belos replacing AztecOO, Ifpack2 replacing Ifpack, MueLu replacing ML, and Tpetra replacing Epetra. The present manuscript introduces these new packages and their capabilities and outlines their interplay with Kokkos.
@@ 2.1. Performance Portability: Kokkos Core
Kokkos Core [^8] For historical reasons the package name is “Kokkos” since it precedes the creation of other Kokkos subprojects such as Kokkos Kernels. is a programming model designed for performance portability across hardware architectures such as CPUs and GPUs. While the development of Kokkos Core has originated as a package of Trilinos, Kokkos has become an independent project in 2015 and is nowadays snapshotted into Trilinos for the pure convenience of dependency management. Two primary insights into achieving performance portability shaped Kokkos’ development: i) abstractions for both parallel execution and data structures are required, and ii) the fundamental programming model must be descriptive rather than prescriptive.
The Kokkos programming model includes execution spaces to specify where an algorithm should run, execution policies to indicate how the algorithm should be run in terms of hardware resources, and execution patterns to designate whether to execute a parallel_for, parallel_reduce, or parallel_scan. The data structure abstractions are built around the concepts of memory spaces to indicate where memory is allocated, memory layouts to specify how data is laid out in memory, and memory traits that enable customizations of how data is accessed. The primary data structure incorporating these concepts is Kokkos::View, a multidimensional array abstraction that inspired the ISO C++23 mdspan class [^9] https://www.open-std.org/jtc1/sc22/wg21/docs/papers/2022/p0009r18.html. Detailed descriptions of Kokkos Core can be found in (Edwards et al., 2014) and (Trott et al., 2022). Development of Kokkos Core itself is carried out in its own GitHub repository [^10] https://github.com/kokkos/kokkos.
@@ 2.2. Performance Portable Kernels: Kokkos Kernels
Kokkos Kernels (Rajamanickam et al., 2021) is part of the Kokkos ecosystem (Trott et al., 2021) and provides node-local implementations of mathematical kernels widely used across packages in Trilinos. As a member of the Kokkos ecosystem, Kokkos Kernels tightly integrates Kokkos features and aims at delivering performance-portable algorithms across major CPU- and GPU-based HPC systems. Due to its node-local nature, Kokkos Kernels does not rely on MPI or other communication libraries, unlike numerous other packages in Trilinos.
@@
The capabilities provided by Kokkos Kernels can be divided into four major categories: 1. BLAS algorithms, 2. sparse linear algebra and preconditioners, 3. graph algorithms, and 4. batched dense and sparse linear algebra (Liegeois et al., 2023; Abdelfattah et al., 2021). The main points of integration of Kokkos Kernels in Trilinos are Tpetra for the dense and sparse linear algebra capabilities, Ifpack2 for the preconditioners and batched algorithms, and the multigrid package MueLu which relies directly and indirectly on BLAS and graph algorithms and specialized implementations of graph coloring/coarsening (Kelley and Rajamanickam, 2022) and fused Jacobi-SpGEMM (generalized sparse matrix-matrix multiply) kernels. Additionally, batched dense factorization algorithms are used in the construction of grid transfers.
@@ 2.3. Tools: Teuchos
Teuchos provides a collection of helper and utility components designed to simplify common programming tasks and to ensure a consistent coding style across Trilinos packages. These tools include memory management classes (Bartlett, 2010) such as smart pointers and arrays, parameter lists for communicating hierarchical lists of parameters between library or application layers, templated wrappers for BLAS and LAPACK, XML parsers, MPI, and other supporting functionality. Teuchos thereby provides a unified “look and feel” across Trilinos packages and helps to avoid common programming mistakes. It offers reusable building blocks for such recurring tasks implemented in a consistent and coherent manner to be used throughout the entire code base.
@@ 2.4. Distributed-Memory Linear Algebra: Tpetra
Tpetra (Baker and Heroux, 2012; Hoemmen, 2015) provides the distributed-memory infrastructure for linear algebra objects such as sparse graphs, sparse matrices, and dense vectors. These linear algebra objects are implemented using MPI-based abstractions for the inter-process communication, and Kokkos-based data structures for the portions of the object owned by each MPI process. Kokkos Kernels is used for performing local, parallel computations within each MPI process. Due to the fundamental nature of MPI distributed data structures, the majority of Trilinos packages depend directly or indirectly on Tpetra.
The objects provided are templated on scalar type (e.g., double, float, or a Sacado or Stokhos type), local index type, global index type, and Kokkos device (pair of execution and memory spaces). The distribution of the linear algebra objects is described by a Map data structure, which defines which elements are assigned to which MPI ranks. Here “element” could be a row or column of a sparse matrix or an entry of a vector. Global indices number elements across an entire distributed object, while local indices number elements only within a rank. The Map encodes the relation between local indices and global indices. An illustration for a vector, that is distributed to two MPI processes, is given in Figure 2. Since memory capacity limits the sizes of local objects, local indices can be safely represented with a more compact type than global indices (e.g., 32-bit int instead of 64-bit long).
[…]
Figure 2. Illustration of a Tpetra::Map for an exemplary vector $\mathbf{v}\in\mathbb{R}^{5}$ distributed to two processes $p\in\{0,1\}$: Entries $v$, $x$ and $z$ are stored on process 0, entries $w$ and $y$ are stored on process 11. Local-to-global mapping of indices is given by $\mathbf{M}_{p}$, where the $i$th element of $\mathbf{M}_{p}$ denotes the unique global index corresponding to the local index $i$ on rank pp. In more complicated scenarios, entries can also be shared between processes.
@@
  • (Multi)Vectors enable to storage of dense vectors or collections of vectors (multivectors) and associated BLAS-1 like kernels (e.g., dot products, norms, scaling, vector addition, pointwise vector multiplication) as well as tall skinny QR (TSQR) factorization for multivectors.
  • CrsGraph implements a sparse graph in compressed row storage (CRS) format. Graphs also include import/export objects for use in halo/boundary exchanges associated with the graph.
  • CrsMatrix and BlockCrsMatrix are sparse matrices in CRS and block compressed sparse row (BSR) formats. Associated format-specific computational kernels include sparse matrix-vector product (SPMV), sparse matrix-matrix multiplication (SPGEMM) and triple-product, sparse matrix-matrix addition, sparse matrix transpose, diagonal extraction, Frobenius norm calculation, and row/column scaling.
@@ 2.5. Abstract Numerical Algorithms: Thyra
The Thyra package provides abstractions for implementing high-level abstract numerical algorithms (ANAs). Thyra is used by other Trilinos packages for implementing ANAs such as solvers for linear and nonlinear equations, bifurcation and stability analysis, optimization, and uncertainty quantification. The abstractions hide concrete linear algebra and solver implementations so that algorithm development is separated from the implementation of the underlying kernels, allowing the same code to be written once and reused across many application codes. For example, one could write a Krylov solver and, without modifications to computational kernels that are hidden behind the abstraction layer, use it with multiple concrete linear algebra implementations such as Tpetra in Trilinos or the vector/matrix objects in PETSc (Balay et al., 2021).
@@
  • Operator Solve Interfaces: The second set of abstractions supplies abstract linear solvers that can be attached to operators. The interfaces include base classes for preconditioners, preconditioner factories, linear operators with a solver, and factories for linear operators with a solver. “Factories” build and update the preconditioner and linear solver objects as needed. The solver interfaces allow one to wrap any concrete linear solver or preconditioner for use with the ANAs. As a concrete implementation of this concept, the Stratimikos package wraps all of the Trilinos preconditioners and linear solvers in Thyra abstractions. Run-time configuration of Stratimikos is driven by Teuchos::ParameterList options. Details on Stratimikos are given in Section 3.8.
  • Nonlinear Interfaces: The final set of abstractions supports nonlinear operations. Thyra provides an interface that wraps nonlinear solvers. Implicit time integration and optimization solvers typically require nonlinear solves and can abstract away the concrete implementation via Thyra. The NOX package in Trilinos (described in Section 4.1) provides a concrete implementation of this interface. The final interface in Thyra is for querying a physics application for common objects required by Newton-based solvers. The Thyra::ModelEvaluator interface allows codes to ask the application to evaluate residuals, Jacobians, parametric sensitivities, Hessians, and post-processed quantities/derivatives of interest. The interface is designed to be stateless (i.e., consecutive calls with the same inputs should produce the same outputs). This design was chosen to avoid pitfalls in complex application codes interacting with general nonlinear algorithms. The design does not preclude its use in stateful applications that have a “memory” between timesteps, but rather forces the explicit use of the stateful information in the interface.